OpenStudy (anonymous):

how to solve a problem like this? Jessica deposits $300 into a savings account that pays an annual interest rate of 2%, compounded twice a year. How much money will Jessica have in her account at the end of one year?

OpenStudy (ash2326):

@Suzy you know compound interest?

OpenStudy (anonymous):

no i just want to know the formula

OpenStudy (ash2326):

Let the deposited amount be P (Principal) r = rate of interest t= time in years n= no. of times the interest is compounded per year n=1 for once a year n=2 for twice a year the final amount is given as \[\large A=P\times (1+\frac{r}{n})^{nt}\] Here we have P=$300 r=2%=(\frac{2}{100}\) n=2 t=1 \[Interest=A-P\]

OpenStudy (ash2326):

Sorry r=2%=\(\frac{2}{100}\)=0.02

OpenStudy (ash2326):

@suzy4321 Can you do it now?

OpenStudy (anonymous):

306

OpenStudy (ash2326):

This is A or interest???

OpenStudy (anonymous):

A meaning answr yes

OpenStudy (ash2326):

A means final amount!!! Did you use Interest= A-P??

OpenStudy (anonymous):

my answer 306.03 is correct

OpenStudy (ash2326):

Yeah it's correct:)